Quantity Total
Cost
Marginal
Cost
Average

Total
Cost

Average
Variable
Cost
0 1,000 - - -
10 1,600 60 160 60
20 2,100 50 105 55
30 2,500 40 83.33 50
40 2,800 30 70 45
50 3,200 40 64 44
60 3,700 50 61.67 45
70 4,300 60 61.43 47.14
80 5,000 70 62.50 50
90 5,800 80 62.44 53.33

Consider the above costs of a purely competitive firm.

Calculate this firm's profit maximizing quantity and profit (or loss) when the market price of the computer software equals $72.

  

quantity = 80; profit = $760

   

quantity = 80; profit = $430

   

quantity = 70; profit = $40

   

quantity = 70; profit = -$170

   

quantity = 50; profit = -$100

   

quantity = 0; profit = -$1,000

   

quantity = 60; profit = $20

   

quantity = 60; profit = -$160

B:

Calculate this firm's profit maximizing quantity and profit (or loss) when the market price of the computer software equals $56.

  

quantity = 60; profit = $220

   

quantity = 50; profit = -$1,050

   

quantity = 30; profit = -$1,210

   

quantity = 70; profit = $40

   

quantity = 60; profit = -$340

   

quantity = 80; profit = $220

   

quantity = 80; profit = -$40

   

quantity = 0; profit = -$1,000
Answer #1

In case of Perfect Competition the firm maximizes profit by equating price to MC.

1. Here price is $72 and at a output of 80 MC = 70

Thus, the firm should produce 80 units.

Profit = $72*80 - $ 5000 = $ 5760 - 5000 = $ 760

First one is correct

Output = 80, profit = $ 760

2. When price is $ 56

Then quantity sold will be 60 units

Profit = $56*60 - 3700 = $ 3360 - 3700 = - $ 340

Quantity = 60, profit = - $ 340

Please contact if having any query thank you.